OpenStudy (anonymous):

Wendell is looking over some data regarding the strength, measured in Pascals (Pa), of some building materials and how the strength relates to the length. The data are represented by the exponential function f(x) = 2^x, where x is the length. Explain how he can convert this equation to a logarithmic function when strength is 8 Pascals.

OpenStudy (anonymous):

I got x = log_2 8

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well you can take the log of both sides and use the log law for powers \[\log(a^b) = b \times \log(a)\] and you should recognise \[8 = 2^3\] so \[\log(2^x) = \log(2^3)\] now use the log law above...

OpenStudy (anonymous):

So the answer is 3 x log(2)?

OpenStudy (campbell_st):

well you need to solve for x so xlog(2) = 3log(2) divide both sides of the equation by log(2) and you get x = 3 the question doesn't need logs... it just needs the skill of identifying 8 as 2^3 hope it helps

OpenStudy (anonymous):

Do I need to solve for x? Doesn't the question just ask to convert the equation to a log function?

OpenStudy (anonymous):

I think the answer must be in the form x = log_b y.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

I agree with your answer :-)

OpenStudy (anonymous):

Thanks. Could you explain why?

OpenStudy (joannablackwelder):

We plug in 8 for f(x) since that is the Pascals.

OpenStudy (joannablackwelder):

Then solve for x using the rules of logs.

OpenStudy (joannablackwelder):

But I agree that it would simplify to 3

OpenStudy (joannablackwelder):

log_2 8 = 3, so it is just a simplification

OpenStudy (anonymous):

Sorry, i meant since 8 = 2^3, does the answer change?

OpenStudy (joannablackwelder):

Sorry, I'm not sure what you mean.

OpenStudy (joannablackwelder):

Well, I was saying that x=log_2 8 which simplifies to x=3

OpenStudy (anonymous):

Are you sure the answer is log_2 8 = 3?

OpenStudy (joannablackwelder):

No

OpenStudy (anonymous):

Okay, well I think that your explanations made a lot of sense, so I will just go with that. Thanks for your time and patience!

OpenStudy (joannablackwelder):

Ok, you're welcome